Find the equation of the line parallel to 10x + 21y = 15 and passing through the point (0,18).
Answer by lwsshak3(11628)

Find the equation of the line parallel to 10x + 21y = 15 and passing through the point (0,18)
**
Given line
10x + 21y = 15
21y=-10x+15
slope m=-10
..
Parallel line
slope=-10
equation:
y=-10x+b
solving for b using coordinates of given point (0,18)
18=0+b
b=18
equation:
y=-10x+18